OpenStudy (anonymous):

The cost, c, of a health club membership is related linearly to its length in months, m. If a 12-month membership costs $260 and an 18-month membership costs $380, which equation can be used to find the cost of a 30-month membership? mc015-1.jpg mc015-2.jpg mc015-3.jpg mc015-4.jpg

OpenStudy (anonymous):

A. C-380=18-12/380-260(30-18) B.C-18=380-260/18-12(380-30) C. C-380=380-260/18-12(30-12) D. C-18=380-260/18-12(30-12)

jimthompson5910 (jim_thompson5910):

can you post a screenshot of the whole thing?